OpenStudy (anonymous):

Simplify. Assume that all variables represent positive real numbers. ^5sqrt(-243x^6y)

jhonyy9 (jhonyy9):

so 5 is the indice of radical ?

OpenStudy (anonymous):

Yes sir its like 5^5 just no number infront of it

jhonyy9 (jhonyy9):

so than you need check wath is the number what multiply 5 time will add to you 243 ?

jhonyy9 (jhonyy9):

like 2*2*2*2*2=32

OpenStudy (anonymous):

Yes but this time the five is inside the sqrt

jhonyy9 (jhonyy9):

yes but this sign indice of radical - so the same like squareroot has indice 2 - or cuberoot has indice 3 - so in this case there is indice 5

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

check 3*3*3*3*3 = ?

OpenStudy (anonymous):

243

jhonyy9 (jhonyy9):

yes right but there is again one minus - so than what sign this ?

OpenStudy (anonymous):

-243

OpenStudy (anonymous):

So its sqrt(-3x^6y)?

jhonyy9 (jhonyy9):

not right 5root(-243x^6y =3x(5root(-xy))

jhonyy9 (jhonyy9):

sqrt sign squareroot so radical indice 2

jhonyy9 (jhonyy9):

do not confuse it never !!!

OpenStudy (anonymous):

so its  ?